# On the existence of unbiased resilient estimators in discrete quantum systems

### v2

@@ B.1 Bhattacharyya bounds for pure states
$Q_{2}=\begin{pmatrix}\langle\psi|\frac{\partial^{2}\rho}{\partial x^{2}}|\psi\rangle&0\\ 2\langle\psi^{\perp}|\frac{\partial^{2}\rho}{\partial x^{2}}|\psi\rangle&0\end{pmatrix},$ (100)
where $\rho=|\psi\rangle\langle\psi|$. From this, one can rewrite the quantum Bhattacharyya matrix (up to the second order) in terms of $Q_{j}$ as
$Q_{BhB}=\begin{pmatrix}\operatorname{Tr}{Q^{\dagger}_{1}Q_{1}}&\frac{1}{2}\left(\operatorname{Tr}{Q^{\dagger}_{1}Q_{2}}+\operatorname{Tr}{Q^{\dagger}_{2}Q_{1}}\right)\\ \frac{1}{2}\left(\operatorname{Tr}{Q^{\dagger}_{1}Q_{2}}+\operatorname{Tr}{Q^{\dagger}_{2}Q_{1}}\right)&\operatorname{Tr}{Q^{\dagger}_{2}Q_{2}}\end{pmatrix}.$ (101)
In this way, one can compute the second orther Bhattacharyya bound as the first element of the inverse of $Q_{BhB}$, i.e. $\Delta_{BhB}=Q^{-1}_{BhB,11}$. In Fig. (3) one can clearly see that the second order Bhattacharyya bound and CR bound differ.
Figure 3: Plot of the variances for different values of $x$. The upper one corresponds to the QBhB bound and the other to CR bound.
L_1 and L_2 formulas: Doing so, one gets
$L_{1}$ $=\left(\begin{array}[]{cc}4\theta(1-2\lambda)\sin\left(2\theta^{2}\right)&4i\theta(2\lambda-1)\cos\left(2\theta^{2}\right)\\ -4i\theta(2\lambda-1)\cos\left(2\theta^{2}\right)&4\theta(2\lambda-1)\sin\left(2\theta^{2}\right)\\ \end{array}\right),$ (104)
$L_{2}$ $=\left(\begin{array}[]{cc}\frac{4(2\lambda-1)\left(\theta^{2}\left(\cos\left(2\theta^{2}\right)-2\lambda+1\right)-(\lambda-1)\lambda\sin\left(2\theta^{2}\right)\right)}{(\lambda-1)\lambda}&\frac{4i(2\lambda-1)\left((\lambda-1)\lambda\cos\left(2\theta^{2}\right)+\theta^{2}\sin\left(2\theta^{2}\right)\right)}{(\lambda-1)\lambda}\\ -\frac{4i(2\lambda-1)\left((\lambda-1)\lambda\cos\left(2\theta^{2}\right)+\theta^{2}\sin\left(2\theta^{2}\right)\right)}{(\lambda-1)\lambda}&\frac{4(2\lambda-1)\left((\lambda-1)\lambda\sin\left(2\theta^{2}\right)-\theta^{2}\left(\cos\left(2\theta^{2}\right)+2\lambda-1\right)\right)}{(\lambda-1)\lambda}\\ \end{array}\right).$ (107)
From this, one can construct the BhBattacharyaa matrix which in this case reads
Appendix C Remarks to Ref. [30]
In this section, we raise some remarks to some claims made in Ref. [30]. Specifically, we present some arguments against the implication that if the matrix $\bm{Q}$ appearing in Eq. (14) is singular, the bound is divergent. Moreover, we also argue against the statement that for pure states all Bhattacharyya bounds give the same value as the CR bound. We make our arguments stronger using simple counterexamples of both propositions. Even though our results are only concerned with the Bhattacharyya bound, we think one could extend them for the other bounds presented in Ref. [30].
C.1 Invertibility of the matrix
In Ref. [30], Gessner and Smerzi argue that if the BhB matrix $\bm{C}$ is singular (non-invertible) the corresponding BhB bound will be divergent. While this may seem true at first inspection, we prove that is not true on a closer look. We now present our objections using the quantum version of the BhB bound, even though the conclusions also hold in the classical case. Moreover, we restrict ourselves to the second order. Let us first invoke the formula which gives the Bhattacharyya bound for the quantum case:
as explained in Ref. [30]. It is easy to see, that if the matrix is singular, one could always obtain zero in the denominator of Eq. (108). However, it does not follow that $\left(\Delta\theta_{\mathrm{est}}\right)_{BhB}^{2}$ will be zero, since it may be the case that the eigenvector with null eigenvalue of $Q_{\mathrm{BhB}}$ may be orthogonal to $\bm{\lambda}$. This would imply that $\left(\Delta\theta_{\mathrm{est}}\right)_{BhB}^{2}$ is zero over zero which is undefined so one cannot say that the variance is unbounded. In addition, we find illuminating to show the discrepancy with the aforementioned claim using a simple example. Consider we have a state given by $\rho=\theta|0\rangle\langle 0|+(1-\theta)|1\rangle\langle 1|$ where $\theta$ is the parameter we would like to estimate. It is clear that this state fulfills $\frac{\partial\rho}{\partial\theta}\neq 0$ and $\frac{\partial^{2}\rho}{\partial\theta^{2}}=0$. Then, $L_{2}=0$ and $L_{1}=1/\theta|0\rangle\langle 0|+1/(\theta-1)|1\rangle\langle 1|$ and the QBhB matrix will read
$\bm{Q}_{BhB^{2}}=\begin{pmatrix}\frac{1}{\theta(1-\theta)}&0\\ 0&0\\ \end{pmatrix},$ (111)
and
$\left(\Delta\theta_{\mathrm{est}}\right)_{BhB}^{2}=\sup_{\bm{a}}\frac{\left(\bm{a}^{\top}\bm{\lambda}\right)^{2}}{\bm{a}^{\top}Q_{\mathrm{BhB}}\bm{a}}=\sup_{\bm{a}}\frac{a^{2}_{1}}{\frac{a^{2}_{1}}{\theta(1-\theta)}}=\frac{1}{\frac{1}{\theta(1-\theta)}},$ (112)
which is essentially the QCRB, since the QFI is equal to $\frac{1}{\theta(1-\theta)}$. So, in this case, even if the QBhB matrix is singular, there exists an estimator that saturates the QBhB even though its value corresponds to the one given by the QCRB.
C.2 Bhattacharyya bounds for pure states
The literature claims that for pure states, Abel quantum bounds and in particular $n^{th}$order QBhB bounds are always equal to QCRB [30]. From our results, this would imply that there is always an estimator satisfying all the unbiasedness conditions. But in fact, it is not true. In the following, we give a simple counter-example and we point out why wrong conclusions are drawn in [30]. Consider the diagonal state $\rho=\lambda|0\rangle\langle 0|+(1-\lambda)|1\rangle\langle 1|$, where in the limit $\lambda\rightarrow 0$ it corresponds to a pure state. The state evolves under the unitary $U_{\theta}=e^{-i\theta\sigma_{x}}$. For this model the first and the third derivatives are equal, $\partial_{\theta}\rho_{\theta}=\partial^{3}_{\theta}\rho_{\theta}$ and
$\bm{Q}_{BhB^{3}}=\left(\begin{array}[]{ccc}4(1-2\lambda)^{2}&0&-16(1-2\lambda)^{2}\\ 0&-\frac{4(1-2\lambda)^{2}}{(\lambda-1)\lambda}&0\\ -16(1-2\lambda)^{2}&0&64(1-2\lambda)^{2}\\ \end{array}\right).$ (116)
Note that the first and the third rows are proportional. The bound in $\lambda\rightarrow 0$ (pure state), is given by
$\left(\Delta\theta_{\mathrm{est}}\right)_{BhB}^{2}=\lim_{\lambda\rightarrow 0}\sup_{\bm{a}}\frac{\left(\bm{a}^{\top}\bm{\lambda}\right)^{2}}{\bm{a}^{\top}\bm{Q}_{\mathrm{BhB}}\bm{a}}.$ (117)
The supreme is achieved for $\bm{a}=(a_{1},0,\frac{a_{1}}{4})$ since it is an eigenvector of $\bm{Q}_{\mathrm{BhB}}$ with eigenvalue 00. In that case one has
$\left(\Delta\theta_{\mathrm{est}}\right)_{BhB}^{2}=\lim_{\lambda\rightarrow 0}\sup_{a_{1}}\frac{a_{1}}{0}=\lim_{\lambda\rightarrow 0}\infty=\infty$ (118)
In this example the BhB bound diverges while the quantum Fisher is finite –and hence the QCRB–, showing that for pure states the quantum Abel bounds and in particular the QBhB bounds can be different than the QCRB and one may be interested in computing them. The key is to consider the limit $\lambda\rightarrow 0$ after the supremum in Eq. (108). In the literature, the limit was taken before the supremum and it was observed that the matrix $\bm{Q}$ had some divergent elements that would not contribute to the bound.

### v3

@@ Appendix B Example of the quantum case
$\bm{C}_{Bh}=\left(\begin{array}[]{cc}\frac{64\theta^{6}(1-2\lambda)^{2}}{4\theta^{4}-\lambda^{2}+\lambda}&0\\ 0&0\\ \end{array}\right)$ (43)
where we have supposed that we are doing the optimal measurement for the point $\theta$, that is, $\phi\to\theta$. Note that $C_{\text{CRB}}$ is not invertible, using (14) it can be noted that the bound is divergent and that there is no estimator satisfying the conditions for 2ndorder BhB. The Fisher information $F_{i}^{\text{CRB}}=16\theta^{2}(1-2\lambda)^{2}$ coincide with QFI, see (41). On the other hand, although $\bm{C}_{Bh}$ is not invertible, if we resort to the maximization (14), we obtain a finite BhB that coincides with the Cramér-Rao bound obtained for $P_{\theta}^{\text{Bh}}$. Thus, we have
$\text{Bh}_{2}=\text{CRB}=\frac{4\theta^{4}-\lambda^{2}+\lambda}{64\theta^{6}(1-2\lambda)^{2}}.$ (44)
The QBh2 derived from (41) coincides with the result obtained in (44). The key insight is that a measurement based on the eigenstates of $(Q^{-1}\bm{\lambda})\bm{L}$ produces a probability distribution with a null second derivative. When the second derivative vanishes, the estimator that saturates the CRB automatically satisfies the conditions for Bh2. This explains why Bh2 and the CRB are identical for the probability distribution $P_{\theta}^{\text{Bh}}$ and why in the quantum case one can obtain more nontrivial higher orders in the hierarchy of bounds.